ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 223, 2018

Barto, L., J. Opršal and M. Pinsker The wonderland of reflections

363

Bergdall, J. and P. Chojecki An adjunction formula for the Emerton–Jacquet functor

1

Bufetov, A. I. and B. Solomyak The Hölder property for the spectrum of translation flows in genus two

205

Cahn, P., F. Fanoni and B. Petri Mapping class group orbits of curves with self-intersections

53

Carmon, D. Evenly divisible rational approximations of quadratic irrationalities

441

Chebotar, M., W.-F. Ke, P.-H. Lee and E. R. Puczyłowski On polynomial rings over nil rings in several variables and the central closure of prime nil rings

309

Chojecki, P. See J. Bergdall Dufloux, L. Projections of Patterson–Sullivan measures and the di-

chotomy of Mohammadi–Oh

399

Fanoni, F. See P. Cahn Forester, M., R. Tang and J. Tao Veech surfaces and simple

closed curves

323

Hall, J. I., Y. Segev and S. Shpectorov Miyamoto involutions in axial algebras of Jordan type half

261

Ke, W.-F. See M. Chebotar Koshlukov, P. E. See F. Y. Yasumura Kupferman, R., C. Maor and R. Rosenthal Non-metricity in the

continuum limit of randomly-distributed point defects

75

Lee, P.-H. See M. Chebotar Maor, C. See R. Kupferman Oppenheim, I. Angle criteria for uniform convergence of averaged pro-

jections and cyclic or random products of projections

343

Opršal, J. See L. Barto Petri, B. See P. Cahn Pinsker, M. See L. Barto Puczyłowski, E. R. See M. Chebotar Rahm, R. and S. Spencer Entropy Bumps and another sufficient

condition for the two-weight boundedness of sparse operators

197

Read, C. J. Banach spaces with no proximinal subspaces of codimension 2

493

Rosenthal, R. See R. Kupferman Schippers, E. Conformal invariants associated with quadratic differ-

entials

449

Segev, Y. See J. I. Hall Shpectorov, S. See J. I. Hall Solomyak, B. See A. I. Bufetov Spencer, S. See R. Rahm Stevenson, B. A Quasi-isometric embedding into the group of

Hamiltonian diffeomorphisms with Hofer’s metric

141

Tang, R. See M. Forester Tao, J. See M. Forester Yasumura, F. Y. and P. E. Koshlukov Asymptotics of graded

codimension of upper triangular matrices 423